Hi can someone show me how to simplify these please??
a.) [(1/x)+1]/(1/x) =
b.) [x/(x+y)]/x =
c.) 3x/(3x+y) =
d.) [1/(x+h)+(1/x)]/x =
i dont know that much about simplifying yet can someone please show me how to simplify these? thank!!

Answer by Nate(3500)

a.) [(1/x)+1]/(1/x) = [(1/x)+1](x) = 1 + x
b.) [x/(x+y)]/x = b.) [x/x(x+y)] = b.) 1/(x+y)
c.) 3x/(3x+y) = already simplified
d.) [1/(x+h)+(1/x)]/x = d.) [1/(x^2+hx)+(x/x)] = 1/(x^2 + hx + 1)